OpenStudy (anonymous):

Tom works for a manufacturing plant and makes $7.00 per hour. Last week he worked 49 hours. For the hours above 40, he receives time-and-a-half. What did Tom earn last week?

OpenStudy (anonymous):

Possible answers A) $311.50 B) $343.00 C) $374.50 D) $514.50

OpenStudy (anonymous):

Are you saying for the 9 hours he worked over the 40 hours, he gets 1.5x more?

OpenStudy (anonymous):

If so, which I'm sure, the answer is then C.

OpenStudy (anonymous):

40 * 7= $280.00 $10.50 * 9 = $94.50 $280.00 +$94.50 -------- $374.50

OpenStudy (anonymous):

$7 * 40 =280 $7 * 1.5$ because he worked over time $10.5 10.5*9=94$ 94$ + 280$ = 374.5$ choice C is right loujoelou you are right :)